help.. I got this wrong,and im totally lost..
state each of the following systems by using either addition or substitution.If a unique solution does not exist,state whether the system is dependent or inconsistent.




10x+ 2y =7
y= -5x+ 3

Answer by Earlsdon(6294)

Solve the system of equations:
1) 10x+2y = 7
2) y = -5x+3
You can use "substitution" to solve this problem. Simply substitute equation 2) into equation 1) and solve for x.
1a) 10x+2(-5x+3) = 7 Simplify and solve for x.
1a) 10x-10x+6 = 7 Combining the x-terms, the x's disappear.
1a) 6 = 7 Never True!
This system of equations represent two parallel lines.
Since there are no ordered pairs (x, y) that satisfy both equations, the system is "inconsistent"